4/7 is the result when 1/2 and x are added
Answer by stanbon(75887)

4/7 is the result when 1/2 and x are added
-----------------
Equation::
(1/2) + x = 4/7
------
Subtract (1/2) from both sides:
x = (4/7)-(1/2)
-----
x = (8-7)/14
----
x = 1/14
---------------
Cheers,
Stan H.
----------